OpenStudy (anonymous):

Name the property the equation illustrates. 9 + (–9) = 0

OpenStudy (anonymous):

http://www.purplemath.com/modules/numbprop.htm

OpenStudy (anonymous):

Commutative Property a + b = b + a

OpenStudy (anonymous):

Associative Property a + (b + c) = (a + b) + c

OpenStudy (anonymous):

Distributive Property a(b + c) = ab + ac